OpenStudy (anonymous):

Solve the exponential equation: 4^4x-2=1/32 =2^8x-4 = 2^8x2^-4 <--- why 2^8x2^-4?

OpenStudy (anonymous):

your job is to see that \[2^2=4\] and also that \[2^{-5}=\frac{1}{32}\] so you can write both of these with the same base

OpenStudy (anonymous):

ok, but why on my solutions is it listed as \[2^{8x}2^{-4}\] I guess i fail to see where they get the 2nd 2 from

OpenStudy (anonymous):

I mean as a step in the problem

OpenStudy (anonymous):

and therefore \[4^{4x-2}=(2^2)^{4x-2}=2^{8x-4}\] making \[2^{8x-4}=2^{-5}\] and so \[8x-4=-5\]

OpenStudy (anonymous):

thus x = -1/8

OpenStudy (anonymous):

yes. i made a typo as well

OpenStudy (anonymous):

Okay, thanks, I was only trying to understand my professor's solution key.. I did not understand her step... but apparently it is not needed.

OpenStudy (anonymous):

oh no, what i wrote is right

OpenStudy (anonymous):

well it is true that \[2^{8x-4}=2^{8x}\times 2^{-4}\] but i am not sure why that helps your cause

OpenStudy (anonymous):

Ahh I see now that you type it like that! I look too deep sometimes. I guess she wants us to understand all steps.. but most of the time it confuses me

OpenStudy (anonymous):

\[2^{8x}=2^{-5}\times 2^{4}=2^{-1}\] so \[8x=-1\] seems like way unnecessary work to me. you already know \[8x-4=-5\] why not just solve that?

OpenStudy (anonymous):

I agree. Wish I could have this program running during our quizzes

OpenStudy (anonymous):

and you know, how you have it typed up there.. that is exactly how she has it typed.